OpenStudy (anonymous):

Simplify. \[\frac{ -x^3 }{ 2 } \times \frac{ 14 }{ 3x^4 }\]

rishavraj (rishavraj):

wht u got???

OpenStudy (anonymous):

Nothing. :( I am not sure what to do after this... \[\frac{ -x^3 }{ 2 } \times \frac{ 14 }{ 3x^4 }\]\[\frac{ -x^3 }{ 2 } \times \frac{ 2 \times 7 }{ 3x^4 }\]

rishavraj (rishavraj):